Use the given information to find the number of degrees of​ freedom, the critical values chi Subscript Upper L Superscript 2 and chi Subscript Upper R Superscript 2​, and the confidence interval estimate of sigma. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Nicotine in menthol cigarettes 95​% ​confidence; n=23​, s=0.29 mg.
Answer #1

Solution :

Given that,

s = 0.29

s2 = 0.0841

n = 23

df = 22

\chi2L = \chi 2\alpha/2,df = 36.781

\chi2R = \chi 21 - \alpha /2,df = 10.982

The 95% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha /2

  \sqrt22 * 0.0841 / 36.781 < \sigma < \sqrt 22 * 0.0841 * 10.982

0.2243 < \sigma < 0.4105

(0.2243 , 0.4105)